THE OPTICAL SPECTRUM OF THE SILICON TERMINATED CARBON CHAINS SiCnH

\underline{D. L. KOKKIN}, IRAP; Université de Toulouse, UPS; CNRS; 9 Av. colonel Roche, BP 44346, F-31028 Toulouse cedex 4, France and Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138; N. J. REILLY, and M. C. McCARTHY, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138; R. C. FORTENBERRY, and T. D. CRAWFORD, Department of Chemistry, Virginia Tech, Blacksburg, VA 24061.

The gas phase optical spectra of the silicon terminated carbon chains, SiC_{n}H (n=3-5) formed in a silane acetylene discharge, have been investigated by R2C2PI and LIF/DF and will be reported here for the first time. Complementary to the experimental work, a theoretical investigation was undertaken with coupled cluster methods to garner a comprehensive understanding of the molecular structures and electronic properties of these systems. For the linear chains where there is an odd number of carbon atoms (SiC_{3}H and SiC_{5}H), the observed transitions are primarily from a ^{2}Π ground state to a ^{2}Σ state, but as in the case of isovalent carbon chains there are some Herzberg-Teller active modes from an excited ^{2}Π state. While a strong Π − Π transition is predicted for SiC_{4}H, the spectrum is dominated by relatively dark sigma state which is vibronically coupled to the bright ^{2}Π state. In contrast to the odd carbon chains, which exhibit relatively sharp spectral features and lifetimes in the 10-100 ns regime, SiC_{4}H shows broadened spectral features consistent with a ca. 10 ps lifetime, and a subsequent long-lived decay (> 30 microseconds) which we tentatively interpret in terms of mixing with a nearby quartet state arising from the same electronic configuration, a process unavailable for the odd chains.